solve the system:
x+y=4
x+z=4
y+z=4

Answer by mducky2(62)

Let's label the equations:
Equation A: x+y=4
Equation B: x+z=4
Equation C: y+z=4

If we subtract A minus B:
(x+y) - (x+z) = 4 - 4
x + y - x - z = 0
y - z = 0
y = z

If we subtract A minus C:
(x+y) - (y+z) = 4 - 4
x + y - y - z = 0
x - z = 0
x = z

Since y = z and x = z:
y = x = z

Plugging back into Equation A:
x + x = 4
2x = 4
x = 2

Since y = x = z:
y = 2
z = 2

Therefore, y, x, and z all equal 2.